The demand and supply schedules for television (TV) sets in Venezuela, a “small” nation that is unable to affect world prices are given as:

Qd= 900-2P

Qs= -200+2P

Suppose Venezuela imports TV sets at a price of $150 each. Under free trade, how many sets does Venezuela produce? _______ Blank 1

How many sets does Venezuela consume? _______ Blank 2

How many sets does Venezuela import? _______ Blank 3 Determine Venezuela's consumer surplus _______ Blank 4 and producer surplus _______ Blank 5.

Assume that Venezuela imposes a quota that limits imports to 300 TV sets. The quota-induced price increase is _______ Blank 6 and the resulting decrease in consumer surplus is _______ Blank 7.

What is the quota's redistributive effect? _______ Blank 8

What is the quota's consumption effect? _______ Blank 9

What is the quota's protective effect? _______ Blank 10

What is the quota's revenue effects? _______ Blank 11

What is the quota's overall welfare loss to Venezuela as a result of the quota? _______ Blank 12
Answer #1

a) Venezuela would produce: -200+2*150 = 100

b) Venezuela would consume: 900-2*150 = 600

c) imports = 600-100 = 500 units

d) Maximum reservation price would be 900/2 = 450

CS = 0.5*600*(450-150) = 90000